Article #6834: Unscheduled Globus service outage
The Globus connection began experiencing issues on Data Depot endpoint around 11:00am EDT. Engineers are currently diagnosing the issue and are workin...
The Globus connection began experiencing issues on Data Depot endpoint around 11:00am EDT. Engineers are currently diagnosing the issue and are workin...
The Bell cluster began experiencing issues with the /scratch filesystem at around 5pm today. Engineers are currently diagnosing the issue and are work...
The Scholar cluster began experiencing issues with adding new students from the courses onto the cluster. Engineers are currently diagnosing the issue...
The Negishi cluster began experiencing issues with SLURM accounting around 8:00am EDT. Engineers are currently diagnosing the issue and are working to...
The Scholar custer down from approximately 4:30PM onwards on Wednesday, March 27, 2024. Engineers have solved the issue and Scholar should return to s...
Clusters Bell, Geddes and Gilbreth experiencing outages as of 4PM Tuesday March 12th. Engineers are working to resolve the problem. Access to these cl...
The Negishi cluster began experiencing issues around 2PM, and engineers isolated the problem with Negishi. In order to resolve the issue, the schedule...
The Anvil cluster began experiencing issues with Slurm Scheduling this past week. Engineers are currently diagnosing the root cause and are working to...
The Geddes cluster began experiencing networking issues around 9:00am EST. Engineers have identified the issue and are working to bring Geddes back to...
Gilbreth is experiencing scheduling issues and jobs have been paused while RCAC works to resolve this issue. Running jobs have also been impacted, so...
Fortress began experiencing issues with its tape library around 5:00PM. Engineers are currently diagnosing the issue and are working to identify a fix...
Multiple clusters have been powered off in MATH G109 datacenter due to a water issue in the building. Affected systems are Bell, Brown, Geddes, Gilbre...
At about noon today (Tuesday 12 September), we discovered an issue with the scheduler database related to the power outage last Sunday. Scheduling o...
Anvil is experiencing more issues related to the power outage yesterday in the Purdue Data Center. Users are currently unable to login via any method,...
Update: As of 3:45pm, the Bell cluster has returned to production status. Scheduling is still paused on the Negishi cluster, and we will have an updat...
Users of Data Depot on RCAC clusters are currently experiencing significant performance degradation. The symptoms manifest as delays in listing or ac...
Open OnDemand services for the Hammer cluster are currently offline. Engineers are investigating a boot disk failure on the server that hosts the gat...
The Hammer cluster began experiencing issues with the Slurm scheduler around 5:00am, Thursday, July 6th. The Slurm scheduler is non-responsive, as a r...
The Geddes cluster began experiencing issues overnight. Engineers are currently diagnosing the issue and are working to identify a fix. Workloads will...
The Data Depot began experiencing issues with its network drive mapping capability around 1:30pm EDT. The symptoms manifest as users being unable to...